Hi I'm struggling with my maths equations. First question is:
Answer these equations by using the substitution method.
y=4x-7
3x+2y=30
Would really appreciate any help.

Answer by Edwin McCravy(20054)

y=4x-7
 3x+2y=30

Take the second equation:

 3x+2y=30

Replace the letter "y" by the right side of the first equation which is
4x-7 placed within parentheses like this: (4x-7)

 3x+2(4x-7)=30

Now use the distributive principle
to remove the parentheses.

 3x+8x-14=30

Combine the like terms 3x and 8x
and get 11x

   11x-14=30

Add 14 to both sides

      11x=44

Divide both sides by 11

        x=4

That's the answer for x,

But you also need the value for y.

So take the first equation

y=4x-7

And replace x by (4)

y=4(4)-7
y=16-7
y=9

So the solution is: x=4 and y=9

Edwin